Grade 11Mechanics

a 10 kg boy standing in a 40 kg boat floating on water is 20 metre from the shore of the river . if he moves 8 metre on the boat towards the shore , then how far is he now from the shore ? assume no friction between boat and water

Aashray Vinayaka
10 Years agoGrade 11

2 Answers

vinay
10 Years ago
To keep center of mass at rest as to net exrernal force is zero so the boy would move 8-x displacement and boat would move x displacement on the opposite direction so
 10*(8-x)=40*(x) so, x=1.6m
8-1.6 = 6.4m
so,
distance from boat = 20-6.4=13.6m
 
Harshil
7 Years ago
10 (8)=(40+10)×y
y=1.6
distance of boy=20-8+1.6
                            =13.6
 
First the boy run 8m towards shore but the boat moves back by some distance y.so the displacment will not be 12 but 12+y